# MAD Speech: Measures of Acoustic Diversity of Speech

### v1

@@ 8.4 Acoustic diversity in Text-to-Speech systems
Bark TTS is a generative text-to-audio system that is based on a decoder-only Transformer model that operates on discretized speech representations, akin to AudioLM Borsos et al. (2023a) and VALL-E Wang et al. (2023). Bark TTS synthesizes utterances in a random voice and acoustic conditions.
@@
StyleTTS 2 combines a style diffusion model with speech with WavLM-based adversarial training Chen et al. (2022a) for an end-to-end training of a multi-component system. We use a checkpoint trained on LibriTTS Zen et al. (2019). The samples are conditioned on a female voice, but randomize timbre and prosody. [^1] https://colab.../Inference_LibriTTS.ipynb
@@
Tortoise TTS combines auto-regressive and diffusion-based components Betker (2023). For training, it pooled LibriTTS Zen et al. (2019), HiFiTTS Bakhturina et al. (2021), and public data sourced from audiobooks and podcasts.
@@
FastSpeech 2 consists of a Transformer-based encoder that embeds phoneme-based text representation, a variance adapter that predicts pitch, energy, and duration values, and a non-autoregressive decoder. We use a publicly available re-implementation trained on LJSpeech Ito and Johnson (2017). [^2] https://github.com/ming024/FastSpeech2
As some of these systems are deterministic, we adopt a different setup and synthesize all transcripts in shuffled LibriTTS test-clean and calculate the average acoustic diversity across batches of audio of size 128. We report the results in Table 12.
Voice Gender Emotion Accent Back. Noise
Bark TTS 39.41 0.90 8.29 8.30 3.84
Tortoise TTS 30.94 0.92 8.73 8.17 3.17
StyleTTS 2 31.54 0.37 7.73 6.81 2.60
FastSpeech 2 19.33 0.29 6.74 5.63 2.42
We see that Bark and Tortoise TTS are the most diverse across all facets, likely due to training on more diverse internet data. Interestingly, Tortoise TTS tends to be more diverse that Bark in emotions while less or equally diverse in other facets. StyleTTS 2 is considerably less diverse in the gender facet — just as one would expect, given that the generation is restricted to female voices. FastSpeech 2 is the least acoustically diverse due to a combination of (a) training on a very homogeneous dataset, and (b) having a mostly non-autoregressive architecture.
All in all, the comparisons we carried out in this Section show that previously proposed changes lead to non-trivial and sometimes even non-monotonic changes in per-facet acoustic diversity. Equally, the sampling temperature, a frequently tuned hyperparameter, also affects the resulting acoustic diversity. We also found that MAD Speech can be useful in detecting acoustic diversity trends in off-the-shelf TTS systems. Overall, we believe those findings highlight the necessity for including acoustic diversity metrics, such as MAD Speech, in the standard evaluations.
9 Limitations and Broader Impact
By building MAD Speech, we make it possible to holistically assess progress in generative speech and reveal possible hidden biases in the acoustic scenes. Yet, our metrics are machine-learned models themselves and are limited by the nature of the data we used to build them. Specifically, the public datasets we use are limited to English language. The labels in the data can be noisy and only represent a limited reflection of the world.
10 Conclusions & Future work
We introduced MAD Speech, a set of metrics for evaluating different aspects of acoustic diversity in speech. MAD Speech metrics are calculated in two steps: (1) a set of utterances is mapped to an embedding space, (2) acoustic diversity is estimated by comparing the resulting embeddings with an aggregation function (cosine dissimilarity or Vendi Score). Calculating embeddings is done by firstly applying a general-purpose representation model with a subsequent metric-specific projection. We focus on five facets of acoustic diversity: voices, gender, emotion, accents, and background noise.
Finally, we highlighted a necessity of controlling acoustic diversity when developing new models and approaches. To this end, we took some recent modelling improvements and demonstrated that they, in fact, affect acoustic diversity of speech. Moreover, we show that even such a standard hyperparameter as sampling temperature has effect on speech diversity. We hope that our metrics would provide a useful tool for a holistic evaluation of developments in the field of speech generation.
Extending MAD Speech to cover more languages and facets is a promising directions of future work.

### v2

@@ Appendix E Acoustic diversity in Text-to-Speech systems - Model details
Bark TTS [^3] https://github.com/suno-ai/bark is a generative text-to-audio system that is based on a decoder-only Transformer model that operates on discretized speech representations, akin to AudioLM Borsos et al. (2023a) and VALL-E Wang et al. (2023). Bark TTS synthesizes utterances in a random voice and acoustic conditions.
@@
StyleTTS 2 (Li et al., 2024) combines a style diffusion model with speech with WavLM-based adversarial training Chen et al. (2022a) for an end-to-end training of a multi-component system. We use a checkpoint trained on LibriTTS Zen et al. (2019). The samples are conditioned on a female voice, but randomize timbre and prosody. [^4] https://colab.../Inference_LibriTTS.ipynb
@@
Tortoise TTS (Betker, 2023a) combines auto-regressive and diffusion-based components Betker (2023b). For training, it pooled LibriTTS Zen et al. (2019), HiFiTTS Bakhturina et al. (2021), and public data sourced from audiobooks and podcasts.
@@
FastSpeech 2 (Ren et al., 2020) consists of a Transformer-based encoder that embeds phoneme-based text representation, a variance adapter that predicts pitch, energy, and duration values, and a non-autoregressive decoder. We use a publicly available re-implementation trained on LJSpeech Ito and Johnson (2017). [^5] https://github.com/ming024/FastSpeech2